OpenStudy (anonymous):

integrate sin(2x)/sin(x). The Answer is 2sin(x)+c why and how?

OpenStudy (anonymous):

The farthest I have gotten is to turn it into sin(2x)*sec

OpenStudy (anonymous):

sin(2x)/sin(x) sin(2x) = sinx (x + x) = sinx cos x + sin x cos x = 2 sinx cosx 2 sinx cosx / sin x 2 cos x Integral of cos x = sin x 2 * sin x

OpenStudy (anonymous):

The answer is 2sin(x)

OpenStudy (anonymous):

How'd you turn sin(2x) into sinx*cosx+sinx*cosx

OpenStudy (anonymous):

Is it just an identity?

OpenStudy (anonymous):

sin ( a + b) = sin a cos b + sin b cos a

OpenStudy (anonymous):

Is it possible to do this by parts?

OpenStudy (anonymous):

I don't know, perhaps, but I don't see why you'd want that. It's so much simpler to just cancel sinx...

OpenStudy (anonymous):

Fair enough.